# Comparing invariant-mass spectroscopy of 8B with ab initio predictions

### v1

@@ Abstract
Levels in 8B have been investigated experimentally using the invariant-mass technique and compared to ab initio calculations. Data sets obtained using =69-MeV 9C and 13O beams on a Be target have been further analyzed to extend the level scheme of 8B for 10 MeV. New levels were observed in the 2+6Li, +3He+, and the +7Be+ exit channels. Momentum correlations between the decay fragments were also investigated in order to deduce the decay pathways and whether the decays are prompt or sequential. This nucleus and its mirror were also investigated in the ab initio symmetry-adapted no-core shell model. Correspondence between the newly observed and predicted levels were made based on the level energy and the decay modes. For positive-parity levels with 3, all predicted levels can be connected to an experimental counterpart (as least tentatively) for 8.4 MeV.
@@ I INTRODUCTION
Light nuclei like 8B and its mirror 8Li are ideal testing grounds for ab initio nuclear approaches as the calculations are computationally tractable. In these approaches, the nuclear force is modeled in the framework of the chiral effective field theory (EFT) (e.g., [Bedaque:2002, Epelbaum:2002, Etem:2003, Epelbaum:2006, Ekstrom:2013, Epelbaum:2014sza, RevModPhys.92.025004]). In particular, chiral potentials start from the degrees of freedom relevant to low-energy nuclear structure, that is, nucleons and pions, and account for the symmetry and symmetry-breaking patterns of the underlying theory of quantum chromodynamics. Understanding low-lying states in these mirror nuclei is crucial for studies of isospin breaking effects, refining nuclear structure models, and probing chiral inter-nucleon interactions.
In this work, we expand the available level scheme of 8B using the invariant-mass technique applied to proton knockout reactions with a 9C beam. Some of the resonances observed in these knockout reactions, have also been verified using projectile fragmentation reactions with a 13O beam. In addition, we confront calculations in the ab initio symmetry-adapted no-core shell model (SA-NCSM) [Launey:2020, DytrychLDRWRBB20] using the NNLOopt chiral potential [Ekstrom:2013] with previously known and newly observed states in 8B. The proton separation energy of 8B is very small (136 keV) and all excited states are resonances, i.e. particle unbound. This study then serves as a test of SA-NCSM calculations in the continuum, which are also benchmarked against energies obtained in the Green’s function Monte Carlo (GFMC) approach [Wiringa:1995, Pieper:2008].
In particular, the SA-NCSM, based on the concept of the no-core shell model (NCSM) [NavratilVB00, BarrettNV13], uses SU(3) symmetry to construct a complete basis, which reduces computational demands while retaining nuclear structure information. An important consequence is that the SA-NCSM can reach ultra-large model spaces and hence, can accommodate large deformation and can couple to continuum degrees of freedom. Furthermore, the SU(3) symmetry-adapted basis of SA-NCSM allows us to study the intrinsic deformation of the low-lying states in 8B, thereby informing about shape evolution and coexistence in 8B and isotopes.
Comparisons of experimental and theoretical level properties allows us to assign spins for the new levels and account for all the predicted 8B levels up to 8.4 MeV. One of the new levels observed is found to undergo a prompt 2 decay. The nature of this 2 emitter is the focus of a companion paper [CharityShortArxiv], where a comparison to the other known =8 2 emitters is made.
@@ II DATA SETS
In this work we make use of three data sets all collected using the Si and CsI(Tl) telescopes from the HiRA apparatus [Wallace:2007] and with secondary beams provided by the Coupled Cyclotron Facility at the National Superconducting Cyclotron Laboratory at Michigan State University. All three experiments utilized a 1-mm-thick Be target and 14 HiRA Si-CsI(Tl) telescopes were arranged in the same configuration 80 cm downstream from the target covering polar angles between 2∘ to 12∘. In two of these data sets, 8B resonances are created from single-proton knockout reactions with an 69-MeV 9C secondary beam. The first of these, labeled as the 9C(1st) data set, was optimized for the detection of 8C decay products (protons and particles) and signals from heavier elements saturated in the Si electronics [Charity:2010, Charity:2011]. Thus this data set is used only for the +3He+ exit channel. The 9C beam for this data set had an intensity of 1.6105 pps and a purity of 65%. In the second 9C data set [9C(2nd)] [Brown:2014], the electronics were upgraded to remove this limitation and the 2+6Li and +7Be exit channels were also examined. The 9C beam for this data set was similar to the first set with an intensity of 1.2105 pps and purity of 52%. However in this experiment, two of the 54 CsI(Tl) detectors did not operate. Unfortunately this included one of the innermost CsI(Tl) detectors, which reduced the efficiency and statistics for the +3He+ exit channel. Thus in the present work we show data from the 9C(1st) data set for this exit channel and from the 9C(2nd) data set for the other two channels. However, the results for the +3He+ exit channel are consistent in the two data sets.
@@
Experimental parallel momentum distributions of 8B resonances from the 9C(2nd) data set have been compared to theoretical models of direct proton knockout in Refs. [Bonaccorso:2018, Charity:2020] and the higher-momenta regions are consistent with this process. At lower momenta there is need for additional yield from presumably more complex multi-step processes. As one expects the ground state of 9C () to be largely -shell, direct knockout should involve the removal of either a or proton producing states of spin 0+, 1+, 2+, or 3+ in 8B. However, given the neighboring 9,10,11N isotopes have -wave ground states [ENSDF, Hooker:2017, Charity:2023b], it is possible there maybe some level of the proton intruder configuration in the ground state of 9C as well. Knockout of an proton would then produce 1- or 2- states.
@@
In order to confirm some of the new resonances observed in the knockout data, we have also examined a data set obtained with an =69.5-MeV beam of 13O of intensity of 3.7105 pps and purity 80% [Webb:2019, Webb:2020, Charity:2021]. Here 8B resonances are created by projectile fragmentation with a net loss of five nucleons from the projectile. Such a reaction may not be as restrictive as to the of the resonances created, but in any case this data set did not yield evidence for any further 8B resonances. However, it did confirm the existence of some of the new resonances found in the one-proton knockout data sets. For the 9C(2nd) data set, the experiment was performed with the CAESAR CsI(Na) -ray detectors [Weissharr:2010] surrounding the target which provided information on coincident rays.
@@ III RESULTS
For the 2+6Li exit channel, one must consider two possibilities for the 6Li fragment. This fragment could be created in its 1+ ground state or in its 0+ second excited state (isobaric analog state) which decays by emitting a 3.562-MeV ray. All other 6Li excited states decay predominantly by particle emission. The excitation energy calculated from the invariant mass assuming the 6Li fragment was in its ground state, i.e. no coincident ray, is designated as . If the 6Li fragment was formed in its 0+ excited state, then the real excitation energy will be larger by the -ray energy (3.562 MeV).
The distributions of obtained from the 9C(2nd) and 13O data sets are displayed in Fig. 1. The narrow peak observed with both data sets at 7.1 MeV is associated with the 0 isobaric analog state in 8B that undergoes prompt 2 emission to the 0+ isobaric analog state in 6Li. As there is an accompanying ray [Brown:2014], its real excitation energy is 10.619 MeV [ENSDF]. The 2 decay of this state from the 9C(2nd) data set has been discussed in Ref. [Brown:2014]. In Fig. 1(a) there is a peak at =8.4 MeV and excess yield is also confirmed in the results from the 13O data set in Fig. 1(a) where a shoulder feature is present.
For the former 9C(2nd) data set, the CAESAR -ray array [Weissharr:2010] was deployed along with HiRA allowing one to determine whether these 2+6Li invariant peaks are associated with the 1+ ground or 0+ isobaric analog state in 6Li. The full-absorption and first-escape peak associated with this ray are clearly seen in the coincidence -ray-energy distribution shown in Fig. 2(a). Using the gate around these peaks, the coincidence spectrum is shown as the data points in Fig. 2(b) which can be compared to the ungated (blue histogram) version. The 7.1-MeV peak is present in the gated spectrum and thus is coincident with the 3.563-MeV ray. Thus its true excitation energy is 3.563-MeV higher. The 8.4-MeV peak can be identified in inclusive spectrum, but is absent from the gated spectrum indicating that this peak is associated with decay to the ground state of 6Li. The two decay paths are illustrated in the level diagram of Fig. 3.
@@
In these fits, the shape of the background is determined using weighted event mixing (Appendix A), but its magnitude is a fit parameter. The centroid and width for the narrow 0 state were fixed to be consistent with the ENSDF value [ENSDF] and a Breit-Wigner intrinsic line shape is assumed. For the other two wider peaks, which may have asymmetric line shapes, we have used profiles consistent with diproton emission in the -matrix approximation [Barker:2001], but for simplicity the two protons are assumed to have zero relative energy. The effects of the detector efficiency and resolution in this and other fits were included using Monte Carlo simulations as in Ref. [Charity:2019]. The fitted parameters for these peaks and their uncertainties are listed in Table 1.
@@
The distribution of 7Be excitation energy obtained from the 3He+ subevents in the 9C(1st) data set is plotted in Fig. 4(a). The 7/2- resonance and a smaller contribution from the 5/2 state are evident in this distribution. Given the strong yield of the 7Be(7/2-) resonance, we have subdivided the events into those inside and outside of the gate which selects this state as an intermediate in the decay. The deduced 8B excitation-energy spectra for these two groups are displayed in Figs. 4(b) and 4(c), respectively. The spectrum associated with the 7/2- intermediate state, Fig. 4(b), is dominated by a peak at 6.1 MeV which was first described in Ref. [Charity:2011] where the correlations measured between the two sequential decay axes were found consistent with either =3± or 4±. As 3- and 4± states cannot be populated by knocking out either a -wave or a -wave proton from 9C, this state must have =3+.
The spectrum generated by vetoing the 7/2- intermediate state, Fig. 4(c), shows the presence of a state at 8.2 MeV and, with much lower yield, another state at 5.4-MeV peak. The red curves in Figs. 4(b) and 4(c) are joint fits to these experimental spectra where the shape of the background contribution is discussed in Appendix B. The intrinsic line shapes are from -matrix theory for sequential decay of isolated resonances [Kalanee:2013]. The 6.1-MeV state, see Fig. 4(b), is taken to decay by -wave proton emission to the 7Be(7/2-) resonance whose intrinsic line shape is taken from Ref. [Charity:2025]. The other states were treated as 3He decays through an 5Li(g.s.) intermediate (see later). The intrinsic line shape of 5Li(g.s.) is taken from the -matrix parameters of [Woods:1988].
@@
It is possible the 8.2 MeV state in the +3He+ channel and the 8.4-MeV peak in the 2+6Li channel are branches on the same state as their centroids and widths are quite similar (see Tables 1 and 2). The lack of an observation of the 8.2-MeV peak in the +3He+ channel from the 13O data set is not significant as this resonance would be difficult to discern above a significantly increased background. Relative to the expected yield for this resonance, the background in the +3He+ channel has increased by a factor of 50 in the 13O data set compared to the 9C data sets.
To infer the nature of the decay mechanism populating the +3He+ channel, we have examined the correlations between the momentum of the three decay products for the gate around the 8.2-MeV peak in Fig. 4. In the fits, the peak yield in this gate represent 20(4)% of the total and thus great care must be made in subtracting the background correlations. The correlations in the background events and their dependence on have been examined extensively, and discussed in Appendix B, for excitation energies above this gate which is almost entirely background in our fits. Correlations for the background events in the gate are then estimated by extrapolation down in energy. The background-subtracted correlations, expressed as the spectra of the relative energies between each fragment type, are plotted in Fig. 5. Note that the gate is applied to the -vetoed spectrum, i.e. the events in Fig. 4(c), and thus the region of the 3He- relative-energy spectrum around the 7Be(7/2-) peak ( gate) is missing. It is clear that our extrapolation of the background correlations is not perfect as the background-subtracted spectra in Figs. 5(a) and 5(c) go slightly negative at high energies. Nevertheless, gross features of the decay mechanism can be inferred.
The distribution of the - relative energy in Fig. 5(a) is dominated by a wide peak at 1.7 MeV consistent with the presence of the broad 5Li ground state. The green curves in all three panels of Fig. 5 are from a simulation of sequential decay via the 8BHe+5Li(g.s.)+3He+ pathway. These curves reproduce the experimental spectra reasonably well, except in the nonphysical regions where the experimental values are negative. An alternative sequential decay scenario could be 8B+7Be(5/2)+3He+, a scenario indicated by the blue curves in Fig. 5. In this second scenario, the 7Be(5/2) intermediate state would show up as a peak in the 3He+ relative energy spectrum shown in Fig. 5(b) at 4.7 MeV. While the experimental data does have a broad peak close to 4.7 MeV, it is smaller and wider than the simulated distribution (blue curve). This simulation makes use of new more-precise measurements of the centroid and width of the 5/2 state [Charity:2025] where the centroid is 300 keV lower than the ENSDF value and the width is roughly half of the ENSDF width.
@@
A presentation of the +7Be decay channel from the 9C(2nd) data set was previously made in Ref. [Brown:2017]. In this section we will extend the analysis of these data and derive limits for the branching ratio of the 8.4-MeV state to this channel. The 1/2- first excited state of 7Be (=0.429 MeV) decays by -ray emission and thus a 8B parent state that decays to the ground and this first excited state would separately contribute to the invariant-mass spectrum offset by the excitation energy of the latter.
@@
The -ray energy spectrum measured in coincidence with the +7Be channel, displayed in Fig. 6(a), shows the presence of 0.429-MeV rays from the decay of the 1/2- first excited state sitting on a significant background. A gate around this peak is used to enhance proton decays to the 1/2- state as compared to the ground state. Inclusive and -gated spectra are shown in Fig. 6(b) and 6(c), respectively. The -gated spectrum shows the enhancement of two peaks, the narrow lowest-energy peak and the broad peak at 4.7 MeV. The former is a small decay branch of the 1+ first excited state of 8B [Brown:2017] whose presence is barely visible in the inclusive spectrum. The main proton decay branch of this level is to the 7Be(g.s.) which corresponds to the peak just higher in energy seen in Fig. 6(c). The 3+ second excited state of 8B is evident by the peak at =2.3 MeV. Assuming -wave proton emission, this state should not decay to the 7Be(1/2-) state and its greatly suppressed presence in the -gated spectrum is presumably due the background contribution in the -ray gate. The broad state at 4.7 MeV, which is only clearly seen in the -gated spectrum, corresponds to a previously unknown 8B state.
@@
The red curves in Figs. 6(b) and 6(c) are from joint fits to the inclusive and -gated spectra. The lineshapes of the 1+ and newly observed state at 4.7 MeV are each treated in the -matrix approximation as isolated resonances with two channels. The magnitude, centroid, and partial width of each channel are varied. These peaks sit upon background distributions (dashed blue curves) generated by the weighted event mixing procedure [Charity:2023]. In addition there is a small fixed contribution (dotted magenta curves) from the 2 decay of the 5.75-MeV level in 9C that sequentially decays through the 8B(3+) intermediate state [Brown:2017]. This component was determined using a fit of the 2+7Be data and simulations of the detector response to fix the contribution when only one of the two protons is detected.
@@
Following the procedure used in previous joint fits of invariant-mass data [Charity:2022], we assume the background in the gate gives rise to a scaled down copy of the inclusive peak yields plus, for the +7Be(1/2-) branches, extra yield fixed by the inclusive yield times the -ray detection efficiency. The +7Be(1/2-) branch of the new level constrains its centroid to =4.666(20) MeV, corresponding to an excitation of 5.095(20) MeV. Its fitted width is =1.15(7) MeV. The branching ratio of this level to 7Be(g.s.) is only restricted to 50% of the 1/2- branch. This state is at the same excitation as a 1- state reported in two studies, MeV [Yamaguchi:2009] and 5.1 MeV [Mitchell:2010], A state of this could be produced by the knockout of an intruder proton in 9C. However, our fitted width is between, and inconsistent with, the reported values of 150(100) keV and 4.6 MeV from [Yamaguchi:2009] and [Mitchell:2010], respectively.
@@
No indication of peaks associated with 1 decay branches of the 8.4-MeV 8B state can be found in these spectra. However, our sensitivity is poor. Such branches will have low detection efficiency as the 8-MeV protons will most often be emitted to laboratory angles outside the acceptance of the apparatus. Only situations where the proton is emitted in the forward or backward directions can they be detected and such events have poor reconstructed invariant-mass resolution [Charity:2019]. Nevertheless, we had added additional peaks in the fits to the distribution associated with 1 decay to the ground and first excited states of 7Be. Correcting for the detection efficiencies of the different branches, the maximum possible yield from the two 1 branches consistent with the experimental data is 2.2 times that found for the 2 branch. In fitting the -gated spectrum alone, the limit for the +7Be(1/2-) branch is 1.3 times the 2 branch. Ignoring the possible contribution from a +3He+ branch, the 2 branching ratio is a least 30%. This reduces to at least 23% if the +3He+ branch is included. Without ambiguity, the 2 branch is a significant deexcitation pathway.

### v2

@@ Abstract
Levels in 8B have been investigated experimentally using the invariant-mass technique and compared with ab initio calculations. Datasets obtained using =69-MeV 9C and 13O beams on a Be target have been further analyzed to extend the level scheme of 8B for 10 MeV. New levels were observed in the 2+6Li, +3He+, and the +7Be+ exit channels. Momentum correlations between the decay fragments were also investigated in order to deduce the decay pathways and whether the decays are prompt or sequential. This nucleus and its mirror were also investigated in the ab initio symmetry-adapted no-core shell model. Correspondence between the newly observed and predicted levels were made based on the level energy and the decay modes. For positive-parity levels with 3, all predicted levels can be connected to an experimental counterpart (as least tentatively) for 8.4 MeV.
@@ I INTRODUCTION
Light nuclei like 8B and its mirror 8Li are ideal testing grounds for ab initio nuclear approaches as the calculations are computationally tractable. In these approaches, the nuclear force is modeled in the framework of the chiral effective-field theory (EFT) (e.g., [6, 27, 25, 28, 24, 26, 30]). In particular, chiral potentials start from the degrees of freedom relevant to low-energy nuclear structure, that is, nucleons and pions, and account for the symmetry and symmetry-breaking patterns of the underlying theory of quantum chromodynamics. Understanding low-lying states in these mirror nuclei is crucial for studies of isospin breaking effects, refining nuclear structure models, and probing chiral inter-nucleon interactions.
In this work, we expand the available level scheme of 8B using the invariant-mass technique applied to proton knockout reactions with a 9C beam. Some of the resonances observed in these knockout reactions, have also been verified using projectile fragmentation reactions with a 13O beam. In addition, we confront calculations in the ab initio symmetry-adapted no-core shell model (SA-NCSM) [34, 22] using the NNLOopt chiral potential [24] with previously known and newly observed states in 8B. The proton separation energy of 8B is very small (136 keV) and all excited states are resonances, i.e. particle unbound. This study then serves as a test of SA-NCSM calculations in the continuum, which are also benchmarked against energies obtained in the Green’s function Monte Carlo (GFMC) approach [55, 46].
In particular, the SA-NCSM, based on the concept of the no-core shell model (NCSM) [42, 5], uses SU(3) symmetry to construct a complete basis, which reduces computational demands while retaining nuclear structure information. An important consequence is that the SA-NCSM can reach ultra-large model spaces and hence, can accommodate large deformation and can couple to continuum degrees of freedom. Furthermore, the SU(3) symmetry-adapted basis of SA-NCSM allows us to study the intrinsic deformation of the low-lying states in 8B, thereby informing about shape evolution and coexistence in 8B and isotopes.
Comparisons of experimental and theoretical level properties allows us to assign spins for the new levels and account for all the predicted 8B levels up to 8.4 MeV. One of the new levels observed is found to undergo a prompt 2 decay. The nature of this 2 emitter is the focus of a companion paper [14], where a comparison to the other known =8 2 emitters is made.
@@ II DATASETS
In this work we make use of three datasets all collected using the Si and CsI(Tl) telescopes from the HiRA apparatus [51] and with secondary beams provided by the Coupled Cyclotron Facility at the National Superconducting Cyclotron Laboratory at Michigan State University. All three experiments utilized a 1-mm-thick Be target and 14 HiRA Si-CsI(Tl) telescopes were arranged in the same configuration 80 cm downstream from the target covering polar angles between 2∘ to 12∘. In two of these datasets, 8B resonances are created from single-proton knockout reactions with an 69-MeV 9C secondary beam. The first of these, labeled the 9C(1st) dataset, was optimized for the detection of 8C decay products (protons and particles) and signals from heavier elements saturated in the Si electronics [13, 12]. Thus this dataset is used only for the +3He+ exit channel. The 9C beam for this dataset had an intensity of 1.6105 pps and a purity of 65%. In the second 9C dataset [9C(2nd)] [8], the electronics were upgraded to remove this limitation and the 2+6Li and +7Be exit channels were also examined. The 9C beam for this dataset was similar to the first set with an intensity of 1.2105 pps and purity of 52%. However in this experiment, two of the 54 CsI(Tl) detectors did not operate. Unfortunately this included one of the innermost CsI(Tl) detectors, which reduced the efficiency and statistics for the +3He+ exit channel. Thus in the present work we show data from the 9C(1st) dataset for this exit channel and from the 9C(2nd) dataset for the other two channels. However, the results for the +3He+ exit channel are consistent in the two datasets.
@@
Experimental parallel momentum distributions of 8B resonances from the 9C(2nd) dataset have been compared with theoretical models of direct proton knockout in Refs. [7, 15] and the higher-momenta regions are consistent with this process. At lower momenta there is need for additional yield from presumably more complex multi-step processes. As one expects the ground state of 9C () to be largely -shell, direct knockout should involve the removal of either a or proton producing states of spin 0+, 1+, 2+, or 3+ in 8B. However, given the neighboring 9,10,11N isotopes have -wave ground states [1, 31, 20], it is possible there maybe some level of the proton intruder configuration in the ground state of 9C as well. Knockout of an proton would then produce 1- or 2- states.
@@
In order to confirm some of the new resonances observed in the knockout data, we have also examined a dataset obtained with an =69.5-MeV beam of 13O of intensity of 3.7105 pps and purity 80% [53, 52, 19]. Here 8B resonances are created by projectile fragmentation with a net loss of five nucleons from the projectile. Such a reaction may not be as restrictive as to the of the resonances created, but in any case this dataset did not yield evidence for any further 8B resonances. However, it did confirm the existence of some of the new resonances found in the one-proton knockout datasets. For the 9C(2nd) dataset, the experiment was performed with the CAESAR CsI(Na) -ray detectors [54] surrounding the target which provided information on coincident rays.
@@ III RESULTS
For the 2+6Li exit channel, one must consider two possibilities for the 6Li fragment. This fragment could be created in its 1+ ground state or in its 0+ second excited state (isobaric analog state) which decays by emitting a 3.562-MeV ray. All other 6Li excited states decay predominantly by particle emission. The excitation energy calculated from the invariant mass assuming the 6Li fragment was in its ground state, i.e. no coincident ray, is designated . If the 6Li fragment was formed in its 0+ excited state, then the real excitation energy will be larger by the -ray energy (3.562 MeV).
The distributions of obtained from the 9C(2nd) and 13O datasets are displayed in Fig. 1. The narrow peak observed in both datasets at 7.1 MeV is associated with the 0 isobaric analog state in 8B that undergoes prompt 2 emission to the 0+ isobaric analog state in 6Li. As there is an accompanying ray [8], its real excitation energy is 10.619 MeV [1]. The 2 decay of this state from the 9C(2nd) dataset has been discussed in Ref. [8]. In Fig. 1(a) there is a peak at =8.4 MeV and excess yield is also confirmed in the results from the 13O dataset in Fig. 1(a) where a shoulder feature is present.
For the former 9C(2nd) dataset, the CAESAR -ray array [54] was deployed along with HiRA allowing one to determine whether these 2+6Li invariant peaks are associated with the 1+ ground or 0+ isobaric analog state in 6Li. The full-absorption and first-escape peak associated with this ray are clearly seen in the coincidence -ray-energy distribution shown in Fig. 2(a). Using the gate around these peaks, the coincidence spectrum is shown as the data points in Fig. 2(b) which can be compared with the ungated (blue histogram) version. The 7.1-MeV peak is present in the gated spectrum and thus is coincident with the 3.563-MeV ray. Thus its true excitation energy is 3.563-MeV higher. The 8.4-MeV peak can be identified in inclusive spectrum, but is absent from the gated spectrum indicating that this peak is associated with decay to the ground state of 6Li. The two decay paths are illustrated in the level diagram of Fig. 3.
@@
In these fits, the shape of the background is determined using weighted event mixing (Appendix A), but its magnitude is a fit parameter. The centroid and width for the narrow 0 state were fixed to be consistent with the ENSDF value [1] and a Breit-Wigner intrinsic line shape is assumed. For the other two wider peaks, which may have asymmetric line shapes, we have used profiles consistent with diproton emission in the -matrix approximation [4], but for simplicity the two protons are assumed to have zero relative energy. The effects of the detector efficiency and resolution in this and other fits were included using Monte Carlo simulations as in Ref. [11]. The fitted parameters for these peaks and their uncertainties are listed in Table 1.
@@
The distribution of 7Be excitation energy obtained from the 3He+ subevents in the 9C(1st) dataset is plotted in Fig. 4(a). The 7/2- resonance and a smaller contribution from the 5/2 state are evident in this distribution. Given the strong yield of the 7Be(7/2-) resonance, we have subdivided the events into those inside and outside of the gate which selects this state as an intermediate in the decay. The deduced 8B excitation-energy spectra for these two groups are displayed in Figs. 4(b) and 4(c), respectively. The spectrum associated with the 7/2- intermediate state, Fig. 4(b), is dominated by a peak at 6.1 MeV which was first described in Ref. [12] where the correlations measured between the two sequential decay axes were found consistent with either =3± or 4±. As 3- and 4± states cannot be populated by knocking out either a -wave or a -wave proton from 9C, this state must have =3+.
The spectrum generated by vetoing the 7/2- intermediate state, Fig. 4(c), shows the presence of a state at 8.2 MeV and, with much lower yield, another state at 5.4-MeV peak. The red curves in Figs. 4(b) and 4(c) are joint fits to these experimental spectra where the shape of the background contribution is discussed in Appendix B. The intrinsic line shapes are from -matrix theory for sequential decay of isolated resonances [2]. The 6.1-MeV state, see Fig. 4(b), is taken to decay by -wave proton emission to the 7Be(7/2-) resonance whose intrinsic line shape is taken from Ref. [18]. The other states were treated as 3He decays through a 5Li(g.s.) intermediate (see later). The intrinsic line shape of 5Li(g.s.) is taken from the -matrix parameters of [56].
@@
It is possible the 8.2 MeV state in the +3He+ channel and the 8.4-MeV peak in the 2+6Li channel are branches on the same state because their centroids and widths are quite similar (see Tables 1 and 2). The lack of an observation of the 8.2-MeV peak in the +3He+ channel from the 13O dataset is not significant because this resonance would be difficult to discern above a significantly increased background. Relative to the expected yield for this resonance, the background in the +3He+ channel has increased by a factor of 50 in the 13O dataset compared with the 9C datasets.
To infer the nature of the decay mechanism populating the +3He+ channel, we have examined the correlations between the momentum of the three decay products for the gate around the 8.2-MeV peak in Fig. 4. In the fits, the peak yield in this gate represent 20(4)% of the total and thus great care must be made in subtracting the background correlations. The correlations in the background events and their dependence on have been examined extensively, and discussed in Appendix B, for excitation energies above this gate which is almost entirely background in our fits. Correlations for the background events in the gate are then estimated by extrapolation down in energy. The background-subtracted correlations, expressed as the spectra of the relative energies between each fragment type, are plotted in Fig. 5. Note that the gate is applied to the -vetoed spectrum, i.e. the events in Fig. 4(c), and thus the region of the 3He- relative-energy spectrum around the 7Be(7/2-) peak ( gate) is missing. It is clear that our extrapolation of the background correlations is not perfect because the background-subtracted spectra in Figs. 5(a) and 5(c) go slightly negative at high energies. Nevertheless, gross features of the decay mechanism can be inferred.
The distribution of the - relative energy in Fig. 5(a) is dominated by a wide peak at 1.7 MeV consistent with the presence of the broad 5Li ground state. The green curves in all three panels of Fig. 5 are from a simulation of sequential decay via the 8BHe+5Li(g.s.)+3He+ pathway. These curves reproduce the experimental spectra reasonably well, except in the nonphysical regions where the experimental values are negative. An alternative sequential decay scenario could be 8B+7Be(5/2)+3He+, a scenario indicated by the blue curves in Fig. 5. In this second scenario, the 7Be(5/2) intermediate state would show up as a peak in the 3He+ relative energy spectrum shown in Fig. 5(b) at 4.7 MeV. While the experimental data do have a broad peak close to 4.7 MeV, it is smaller and wider than the simulated distribution (blue curve). This simulation makes use of new more-precise measurements of the centroid and width of the 5/2 state [18] where the centroid is 300 keV lower than the ENSDF value and the width is roughly half of the ENSDF width.
@@
A presentation of the +7Be decay channel from the 9C(2nd) dataset was previously made in Ref. [9]. In this section we extend the analysis of these data and derive limits for the branching ratio of the 8.4-MeV state to this channel. The 1/2- first excited state of 7Be (=0.429 MeV) decays by -ray emission. Thus a 8B parent state that decays to both the ground and this first excited state would contribute two peaks to the invariant-mass spectrum separated from each other by 0.429 MeV.
@@
The -ray energy spectrum measured in coincidence with the +7Be channel, displayed in Fig. 6(a), shows the presence of 0.429-MeV rays from the decay of the 1/2- first excited state sitting on a significant background. A gate around this peak is used to enhance proton decays to the 1/2- state as compared with the ground state. Inclusive and -gated spectra are shown in Fig. 6(b) and 6(c), respectively. The -gated spectrum shows the enhancement of two peaks, the narrow lowest-energy peak and the broad peak at 4.7 MeV. The former is a small decay branch of the 1+ first excited state of 8B [9] whose presence is barely visible in the inclusive spectrum. The main proton decay branch of this level is to the 7Be(g.s.) which corresponds to the peak just higher in energy seen in Fig. 6(c). The 3+ second excited state of 8B is evident by the peak at =2.3 MeV. Assuming -wave proton emission, this state should not decay to the 7Be(1/2-) state and its greatly suppressed presence in the -gated spectrum is presumably due the background contribution in the -ray gate. The broad state at 4.7 MeV, which is only clearly seen in the -gated spectrum, corresponds to a previously unknown 8B state.
@@
The red curves in Figs. 6(b) and 6(c) are from joint fits to the inclusive and -gated spectra. The lineshapes of the 1+ and newly observed state at 4.7 MeV are each treated in the -matrix approximation as isolated resonances with two channels. The magnitude, centroid, and partial width of each channel are varied. These peaks sit upon background distributions (dashed blue curves) generated by the weighted event mixing procedure [17]. In addition there is a small fixed contribution (dotted magenta curves) from the 2 decay of the 5.75-MeV level in 9C that sequentially decays through the 8B(3+) intermediate state [9]. This component was determined using a fit of the 2+7Be data and simulations of the detector response to fix the contribution when only one of the two protons is detected.
@@
Following the procedure used in previous joint fits of invariant-mass data [16], we assume the background in the gate gives rise to a scaled down copy of the inclusive peak yields plus, for the +7Be(1/2-) branches, extra yield fixed by the inclusive yield times the -ray detection efficiency. The +7Be(1/2-) branch of the new level constrains its centroid to =4.666(20) MeV, corresponding to an excitation of 5.095(20) MeV. Its fitted width is =1.15(7) MeV. The branching ratio of this level to 7Be(g.s.) is only restricted to 50% of the 1/2- branch. This state is at the same excitation as a 1- state reported in two studies, MeV [57] and 5.1 MeV [40], A state of this could be produced by the knockout of a intruder proton in 9C. However, our fitted width is between, and inconsistent with, the reported values of 150(100) keV and 4.6 MeV from [57] and [40], respectively.
@@
No indication of peaks associated with 1 decay branches of the 8.4-MeV 8B state can be found in these spectra. However, our sensitivity is poor. Such branches will have low detection efficiency as the 8-MeV protons will most often be emitted to laboratory angles outside the acceptance of the apparatus. Only situations where the proton is emitted in the forward or backward directions can they be detected and such events have poor reconstructed invariant-mass resolution [11]. Nevertheless, we had added additional peaks in the fits to the distribution associated with 1 decay to the ground and first excited states of 7Be. Correcting for the detection efficiencies of the different branches, the maximum possible yield from the two 1 branches consistent with the experimental data is 2.2 times that found for the 2 branch. In fitting the -gated spectrum alone, the limit for the +7Be(1/2-) branch is 1.3 times the 2 branch. Ignoring the possible contribution from a +3He+ branch, the 2 branching ratio is a least 30%. This reduces to at least 23% if the +3He+ branch is included. Without ambiguity, the 2 branch is a significant deexcitation pathway.